OpenStudy (anonymous):

solve 3 + sqrt z-11= sqrt z+4

OpenStudy (anonymous):

\[3+\sqrt{z}-11=\sqrt{z}+4\] is it like this? or \[3+\sqrt{z-11}=\sqrt{z}+4\] or something else?

OpenStudy (anonymous):

z is a complex thomas?

OpenStudy (anonymous):

i don't know but if it's complex then i'm not familiar with those :D

OpenStudy (anonymous):

z =x+yi and you will take the square of both sides of you write last

OpenStudy (anonymous):

its how you wrote it Thomas

OpenStudy (anonymous):

the second one..

OpenStudy (anonymous):

so is it complex number or not? :D